M.Y University claims that some of their BS accountancy first year students shifted to online classes because of the difficulties they’d encountered, which is normally distributed. A random sample of 200 students revealed that 55 of them shifted.


Compute for the proportion estimate of the parameter.


Compute for the margin of error of the proportion at 90% CL.


Find 95% CI for all the students who shifted.

Interpret the result.


can someone solve this please


1
Expert's answer
2021-05-31T09:41:47-0400

a) The sample proportion is computed as follows, based on the sample size "N=200"

and the number of favorable cases "X=55:"


"\hat{p}=\dfrac{X}{N}=\dfrac{55}{200}=0.275"

b) The critical value for "\alpha=0.1" is "z_{1-\alpha/2}=1.6449."

The margin of error of the proportion at 90% CL


"ME=z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}"

"=1.6449\times\sqrt{\dfrac{0.275(1-0.275)}{200}}"

"\approx 0.051935"

c) The sample proportion is computed as follows, based on the sample size "N=200"

and the number of favorable cases "X=55:"


"\hat{p}=\dfrac{X}{N}=\dfrac{55}{200}=0.275"

The critical value for "\alpha=0.05" is "z_{1-\alpha/2}=1.96."


The corresponding confidence interval is computed as shown below:


"CI(Proportion)"

"=(\hat{p}-z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}}, \hat{p}+z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=\bigg(0.275-1.96\times\sqrt{\dfrac{0.275(1-0.275)}{200}},"

"0.275+1.96\times\sqrt{\dfrac{0.275(1-0.275)}{200}}\bigg)"

"=(0.213116, 0.336884)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.213116<p<0.336884)," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.213116, 0.336884)."